# Dynamical correlation effects in twisted bilayer graphene under strain and lattice relaxation

### v1

@@ Abstract
We study the impact of lattice effects due to heterostrain and relaxation on the correlated electron physics of magic-angle twisted bilayer graphene, by applying dynamical mean-field theory to the topological heavy fermion model. Heterostrain is responsible for splitting the 8-fold degenerate flat bands into two 4-fold degenerate subsets, while relaxation breaks the particle-hole symmetry of the unperturbed THF model. The interplay of dynamical correlation effects and lattice symmetry breaking enables us to satisfactorily reproduce a wide set of experimentally observed features: splitting the flat band degeneracy has observable consequences in the form of a filling-independent maximum in the spectral density away from zero bias, which faithfully reproduces scanning tunneling microscopy and quantum twisting microscopy results alike. We also observe an overall reduction in the size and degeneracy of local moments upon lowering the temperature, in agreement with entropy measurements. The absence of particle-hole symmetry has as a consequence the stronger suppression of local moments on the hole-doped side relatively to the electron-doped side, and ultimately causes the differences in existence and stability of the correlated phases for negative and positive doping. Our results show that even fine-level structures in the experimental data can now be faithfully reproduced and understood.
@@
Flat bands at the Fermi level in magic-angle twisted bilayer graphene [1] result in strong electron correlation and topological effects, giving rise to a rich low-temperature phase diagram including multiple superconducting domes at various fillings within the flat-band manifold [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. A comprehensive understanding of the low-temperature phases, including the precursors to the superconducting state, necessitates a natural description of strong correlation effects in faithful models that include all the important features of the system, accounting for unavoidable perturbations such as strain and lattice relaxation effects. In this work, we apply dynamical mean-field theory (DMFT) to an extension [30] of the topological heavy fermion model (THF) [31] of twisted bilayer graphene (TBLG), that incorporates lattice effects, by which we refer to the atomistic deformation of the graphene honeycomb structure due to lattice relaxation and strain [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 6, 54].
Lattice relaxation follows from the energetics of the emergent moiré pattern, as locally AB-stacked regions are energetically favored compared to the locally AA-stacked regions. This results in both in-plane and out-of-plane deformations of the atomistic graphene structure, causing the AA regions to shrink relative to the AB regions and bringing the two layers closer to each other at the AB regions compared to the AA regions [55, 56, 57, 58, 59, 60, 61, 7, 62, 63, 44, 51, 50, 64, 65]
@@
In addition, TBLG samples are typically reported to have substantial heterostrain ($\epsilon=0.1\%$$0.4\%$), stemming from several sources during sample preparation and handling, with ultralow-strain samples ($\epsilon<0.1\%$) being rare [11, 60, 61, 66, 67, 68]. Typical and ultralow-strain samples are known to feature different correlated insulating states at low temperatures [11], highlighting the need for the modeling of electron correlations in the presence of lattice effects [69, 65, 70, 71, 58, 2, 4, 59, 60, 61, 7, 62, 66, 8, 56, 72, 63, 73, 74, 9, 67, 75, 64, 65, 76, 77, 68].
While both the original Bistritzer-MacDonald (BM) [1] and the THF model [31] ignore these lattice effects, they have been successful in capturing many aspects of the low-temperature flat band physics. Recently, good progress has been achieved through the study of the THF [78, 79, 80] and other TBLG models [81, 82, 83, 84, 85] within the framework of DMFT and QMC. In the context of the THF, which is at its core a generalized periodic Anderson model (PAM) [86], this has proven both natural and fruitful, opening the way to explanations and predictions of many distinct experimental properties such as the onset of ordering, transport properties [87] and response functions including susceptibility and compressibility. In particular, spectral functions that are very different from those of the non-interacting BM model have been predicted [78, 88, 81] and recently observed [89] by quantum twisting microscopy.
However, a set of experimental features have so far eluded a theoretical prediction based on DMFT, in particular
  1. 1.
    The emergence of a filling-independent persistent spectral feature at $V_{b}\sim 10$ meV in the local spectral function [75, 89, 70].
  2. 2.
    The temperature and filling dependence of entropic measurements suggesting a transition from an $8$-fold degenerate local moment to a $4$-fold degenerate local moment when lowering the temperature. [70, 90, 91].
  3. 3.
    The particle-hole asymmetry of positive vs negative fillings measured from the charge neutrality point (CNP): across samples, the charge compressibility varies more strongly on the electron (e)-doped side [91, 90, 92, 93, 75], and superconductivity shows higher stability on the hole (h)-doped side [2, 20, 25]. At very low T, symmetry broken integer filling insulators are more stable on the electron side, a feature well-reproduced by Hartree-Fock studies [94, 95].
Here, we study a generalization of the topological heavy Fermion model [30], incorporating strain and lattice relaxation terms [33, 49, 44]. We employ the DMFT framework developed in [78, 79] to study the effects of correlations in the presence of strain and lattice relaxation in absence of spontaneous symmetry breaking. We show that the three features above, which persist in the symmetric phase, are correctly accounted for by this approach.
@@
Both strain and lattice relaxation effects reduce the symmetries of the system. In the absence of perturbations, in addition to the translational symmetries and $SU(2)$ spin symmetry, the valley-projected BM model possesses $C_{3z},C_{2x},C_{2z},\mathcal{T}$ symmetries and an emergent anti-commuting particle-hole symmetry $P$. Heterostrain generically breaks $C_{3z}$ and $C_{2x}$, while lattice relaxation breaks $P$ [30]. The breaking of $P$ by relaxation causes the upper flat band (per valley, per spin) to become more dispersive than the lower flat bands.
In absence of strain and lattice relaxation, the narrow-band manifold possesses a $U(2)\times U(2)$ symmetry, which in the chiral flat limit [96, 97, 98] becomes $U(4)\times U(4)$. The addition of lattice effects provides a considerable deviation from this scenario, splitting the flat-band manifold into two four-fold degenerate subsets offset by an effective crystal field proportional to the strain parameter [30].
@@
The breaking of some of the lattice and particle-hole symmetries, together with a proper evaluation of electronic correlations, are the key to unlock an accurate description of the previously listed phenomena. We find that typical strain values $\sim$$0.1\%$ induce a crystal-field–like splitting of the flat-band manifold of the order of $10$ meV. This directly results in a $\sim$$10$ meV persistent feature in the local density, that matches the filling-independent peak reported in [75, 89]. Furthermore, our results show that relaxation-induced particle-hole asymmetry in the non-interacting dispersion carries over to the correlated physics, with stronger variations in the charge compressibility on the electron-doped side compared to the hole-doped side.
@@
Interestingly, we find that in typically strained samples, when the system is doped away from the charge neutrality point, correlations are largely frozen in the inactive (away from the Fermi level) flat-band sector, and present in the active (crossing the Fermi level) flat-band sector. This has strong implications for the degeneracy of the local moments, as is observed in the entropy versus filling curve, which progressively vanishes at the CNP as the temperature is lowered below the characteristic scale of the strain splitting, in good agreement with recent experimental data [70].
@@
The THF model accurately maps the BM continuum model for twisted bilayer graphene to a generalized periodic Anderson model [31]. Per moiré unit cell, two localized $f$ modes (per spin and valley) represent the maximally localized Wannier functions, which correspond to $p_{x}\pm ip_{y}$ orbitals localized at the locally AA-stacked regions and capture $\sim 96\%$ of the flat band spectral weight. Four dispersive $c$-modes (per spin and valley) make up most of the two closest remote bands above and below the flat-band manifold and reproduce the correct band representation of the flat bands by virtue of a band inversion at the $\Gamma_{M}$ point. Per spin and valley, the non-interacting Hamiltonian can be written in the form of a $6\times 6$ matrix $H^{\eta}_{HF}(k)$ acting on the spinor $(f^{\eta}_{\alpha},c^{\eta}_{a})$ with $a=1,2,3,4$ and $\alpha=1,2$ and $\eta=K,K^{\prime}$ representing the valley.
Ref. [30] shows that the modifications to the band structure due to strain and lattice relaxation can be incorporated—directly from ab initio continuum studies [33, 49, 44, 41, 99]—into the THF model via first-order perturbation theory, resulting in the perturbed Hamiltonian $H^{\eta}_{HF,\epsilon,\Lambda}$,
@@
where $\delta H^{\eta}_{\epsilon}$, $\delta H^{\eta}_{\Lambda}(\mathbf{k})$ represent the corrections due to strain and relaxation respectively, $\epsilon$ parametrizes the strain, and $\Lambda$ parametrizes the non-local tunneling terms stemming from lattice relaxation. The primary effect of $\delta H^{\eta}_{\Lambda}$ is the introduction of two orbital-dependent chemical potential terms that shift the two pairs of $c$-electrons relative to each other and to the $f$-electrons, thus breaking $P$-symmetry. $\delta H_{\epsilon}^{\eta}$ introduces 6 strain parameters that couple to the strain tensor. In the following, we will use the parametrization described in [30], the details of which are reproduced in section S1 of the supplementary material [100]. Following the typical strain values reported in experimental samples, we use $0.15\%$ uniaxial heterostrain throughout this paper. Fig. 1(a) shows the typical band structure with strain and lattice relaxation included. $\delta H_{\epsilon}^{\eta}$ crucially contains the term $M_{f}\epsilon_{-}\sigma_{y}$ in the $2\times 2$ $f$-sector, where $\epsilon_{-}$ is the anisotropic component of the strain tensor, and $M_{f}=4380$ meV [30] causing the flat-band manifold to split with finite strain. At $0.15\%$ strain (used throughout this paper), the split is about $7$ meV. In the following, we will refer to the split narrow bands as the bonding and anti-bonding flat bands, and to the local eigenfunctions of the strained $f$-subspace Hamiltonian (which are the eigenstates of $\sigma_{y}$) as the $f_{-}$ and $f_{+}$ orbitals.
We solve the interacting problem via charge self-consistent DMFT (see [78] and section S4 of the supplementary material [100] for details), assuming a double-gate screened interaction with an inter-gate distance $\xi=10$ nm and dielectric constant $\epsilon=6$. We treat the $f$-subspace local interactions at all-order, and all other two-body terms at the Hartree level by coupling to the self-consistently adjusted density matrix. We disregard spontaneous symmetry breaking by setting a temperature of $T\geq 11.6K$, above the onset of ordering [78].
Figure 1: (a) Non-interacting dispersion of the unperturbed (dashed) and strained and relaxed (solid) THF models for the $K$ valley. (b) Occupation of the $f_{+}$ and $f_{-}$ orbitals as a function of filling. On the e-doped side, the $f_{+}$ electrons are active while the $f_{-}$ electrons are nearly fully occupied and vice versa for the h-doped side. (c) Local (mBZ-averaged) spectral function, (d) spectral function at the $K_{M}$-point and (e) spectral function at the $\Gamma_{M}$-point for fillings $\nu\in[-4,4]$ with $0.15\%$ uniaxial heterostrain and lattice relaxation. Strain-induced splitting of the flat-band manifold causes a persistent feature at about ++()(-)$10$ meV on the h-(e-)doped side. The black dotted line in the (e) is the $\mu(\nu)$ curve. The dashed vertical lines are drawn at $\pm M_{f}\epsilon_{-}$ which is the magnitude of the strain-induced splitting in the non-interacting model.
.3 Spectral cascades and charge sector freezing
In Fig. 1(c) we show how the interacting spectral function of TBLG evolves with filling when both strain and lattice relaxation are included. These figures can be compared directly with the local density measured via scanning tunneling microscopy (STM) [75, 55, 59, 60, 61, 56, 73, 64, 57] and the quantum twisting microscope [89, 11, 73], as well as the previous DMFT data on unstrained models [78, 81]. Similarly to the unstrained and unrelaxed system, the spectral function features cascade transitions, with side bands forming patterns that repeat every time the number of electrons per moiré unit cell changes by an integer, and the $f$- and $c$-electron occupations reset. However, the spectra from the perturbed model reproduce one important feature in the experimental data that has thus far eluded explanation. Quantum Twisting Microscope (QTM) [89] and STM [75] spectra report the presence of a “persistent” (filling-independent) feature at around $\sim$$10$ meV on the h-doped site and $\sim$10-10 meV on the e-doped side. While this feature was absent in previous unstrained DMFT calculations [78, 81], our spectral functions of the perturbed model reproduce the persistent feature (cf. the dashed vertical lines in Fig. 1(c-e)) and provide a natural explanation for the corresponding excitations.
In Fig. 1(c-e) we show the local spectral function (i.e. averaged over the mini BZ of BLG), as well as the spectral functions at the $K_{M}$ and $\Gamma_{M}$ high-symmetry points. The theoretical data are in good agreement with the QTM results at the corresponding momenta [89]. At the $K_{M}$ point, the spectrum shows a well defined gap around the Fermi energy, which is maximum around the CNP. The local spectral function, however, shows a residual spectral resonance crossing the gap. This is mostly due to the $c$-electron spectral weight at $\Gamma_{M}$, as explained in detail in section S5 of the supplementary material [100]. Since the $c$-electrons are weakly correlated, the associated spectral peaks effectively behave as noninteracting bands, rigidly shifting upon chemical potential variations. This gives rise to the striking correspondence between the central spectral peak and the $\mu(\nu)$ curve, represented by the black dotted line in  Fig. 1(e), which is also observed experimentally in [89].
The breaking of $C_{3z}$ by strain causes the $8$-fold degenerate flat-band manifold in the unperturbed THF to split into two sets of degenerate anti-bonding and bonding $f$-bands. In Fig. 1(b), we follow the occupation of the $f_{+}$ and $f_{-}$ electrons as a function of total filling $\nu$. We find that on the e-doped side, the $f_{-}$ manifold occupation is almost complete and constant, while the $f_{+}$ electrons are active, and vice versa on the h-doped side. This suggests that the persistent feature is related to excitations to the inactive sector. Since this switches at the charge neutrality point, it appears at positive or negative bias depending on whether the system is hole- or electron-doped, and the energy at which it appears depends on the magnitude of strain in the sample.
We confirm this hypothesis by looking at the momentum-resolved spectral function in Fig. 2, showing the momentum-resolved spectral function for $T=11.6K$ at $\nu=-0.8$. The hue represents whether the spectral weight comes from the $f_{-}$ (blue) or $f_{+}$ (red) orbitals. Since this is on the hole-doped side of CNP, the occupation of the $f_{+}$ (red) electrons is nearly zero. Dynamical correlations are unimportant in this sector and the energy of the corresponding spectral weight is set by the strain splitting. This spectral weight is precisely the content of the persistent feature (marked by the dashed horizontal line in  Fig. 1), and at this filling, is composed of $f_{+}$ (red) electrons. The active sector is composed of the $f_{-}$ (blue) electrons, whose occupation changes as a function of filling. This sector exhibits dynamical correlations, developing the Hubbard bands and the zero-bias resonance forming the cascades. At the considered filling, this spectral weight is pinned close to the Fermi level. In agreement with the observations in the QTM experiment [89], this contribution is absent at the $\Gamma_{M}$ point. We arrive at the conclusion that strain causes the flat bands to split (single particle effect) into anti-bonding and bonding subsets, only one of which is active depending on whether the system is electron- or hole-doped. The inactive sector causes the filling-independent persistent feature, while the active sector is responsible for the correlations and cascade resets. Note that the freezing of the $f_{+}$ or $f_{-}$ sector suggests that the strained system could admit a truncated correlated electron model away from the charge neutrality point in the same vein as the as the “partial symmetry breaking approach” [101, 80] when applied to the strain-split basis. This would allow impurity solvers that would otherwise be computationally unfeasible, such as the Numerical Renormalization Group, to be used within a controlled approximation.
@@
Figure 2: Momentum-resolved $f$-projected spectral function with lattice relaxation and $0.15\%$ strain at filling $\nu=-0.8$ projected onto the strain split $f_{+}$ (red) and $f_{-}$ (blue) orbitals. On the h-doped side ($\nu<0$), the persistent feature at around $10$ meV is made up of unoccupied $f_{+}$ spectral weight.
$P$-symmetry breaking and inverse compressibility– As previously discussed, lattice relaxation effects are responsible for particle-hole symmetry breaking in the THF model. To explicitly quantify them, we discuss the relation between the chemical potential and the total filling. Fig. 3(b) shows the inverse charge compressibility, $\partial\mu/\partial\nu$, of the strained and relaxed THF model compared to the unperturbed THF model from  [78] and experimental data [102, 93, 91].
We find that the addition of strain and lattice relaxation terms allows us to capture an important qualitative feature missing in the unperturbed THF model. Multiple experiments [102, 93, 91] across different samples consistently find a strong particle-hole asymmetry: the variations in the charge compressibility are markedly stronger on the e-doped side compared to the h-doped side. The THF model with strain and lattice relaxation reproduces this phenomenon, and it can be understood by considering the primary effect of the lattice relaxation part of the Hamiltonian, which is a relative downward shift of the $f$-electron energies with respect to the charge neutrality point of the $c$ Dirac bands. This is shown schematically in Fig. 3(a) where the bare $f$ dispersion is represented by the red line, shifted down with respect to the touching point of the $c$ bands. As a consequence, there is a higher $c$ density of states available on the hole-doped side compared to the electron-doped side, and therefore there is stronger hybridization on the hole-doped side compared to the electron-doped side. This suppresses the local moment on the hole-doped side relative to the electron-doped side, which results in the softening of the compressibility variations. This is consistent with Hartree-Fock simulations which find that the ordered insulators on the electron-doped side have a larger gap than the ones on the hole-doped side [95]. Note that one might naively expect that since the upper flat band is more dispersive, correlations would be stronger on the hole-doped side. However, the crucial factor affecting the strength of correlations is the hybridization, which in turn correlates positively with the higher $c$ density of states available on the hole-doped side.
Another key improvement with respect to previous simulations regards the position and shape of the inverse compressibility maxima. Though different experimental results show variations in the position and relative size of the peaks for electron doping (positive $\nu$), the behavior for hole doping (negative $\nu$) is remarkably consistent [91, 90, 93, 102], and shows a marked depinning of the inverse compressibility maxima from integer total occupations. While some theoretical simulations using Gutzwiller variational methods in the unperturbed THF have been able to replicate this feature [88], generically, this model (dashed black line in Fig. 3) features extremely narrow compressibility divergencies up to $\nu=-3$, which are moreover almost exactly pinned at integer occupations [78]. The inclusion of strain and relaxation terms immediately leads to a better estimation of the position and size of the compressibility maxima on the hole-doping side (blue solid line). The increased hybridization on the hole-doped side again plays a crucial role in the depinning. The hybridization has a smoothening effect making the $f$-sector less quantum-dot like, lowering the penalty of non-integer occupation. In other words, the system is more metallic on the hole-doped side than on the electron-doped side (see further discussion in section S5 of the SI [100]). Further softening in the experimental data is likely related to twist-angle and filling inhomogeneities of the samples.

### v2

# Interplay between many-body correlations, strain and lattice relaxation in twisted bilayer graphene

@@ Abstract
In twisted bilayer graphene, a unified understanding of the mechanisms governing temperature-dependent electronic spectra and thermodynamic properties remains controversial despite extensive theoretical efforts. Here, we present a comprehensive theoretical framework that quantitatively accounts for scanning tunneling spectroscopy, quantum twisting microscopy, and thermodynamic properties of magic angle twisted bilayer graphene. We demonstrate that the observed behavior arises from the interplay between electron correlations and external symmetry-breaking induced by strain and lattice relaxation. These effects act cooperatively to shape the emergent electronic behavior, leaving characteristic signatures across spectroscopy, compressibility and entropy.
Figure 1: a Sketch of strained twisted bilayer graphene lattice structure. The red arrows represent the moiré lattice vectors, while the shaded area is the strained unit cell. b Non-interacting dispersion of the unperturbed (dashed) and strained and relaxed (solid) THF models for the $K$ valley. c Occupation of the $f_{+}$ and $f_{-}$ orbitals as a function of filling. d Local (mBZ-averaged) spectral function. e Spectral function at the $K_{M}$-point and f at the $\Gamma_{M}$-point for fillings $\nu\in[-4,4]$ with $0.15\%$ uniaxial heterostrain and lattice relaxation. All data are in the symmetric phase and for $T=11.6$ K. The black dotted line in f is the $\mu(\nu)$ curve. The dashed vertical lines are drawn at $\pm M_{f}\epsilon_{-}$ which is the magnitude of the strain-induced splitting in the non-interacting model.
@@
Strain and lattice relaxation are ubiquitous effects in graphene moiré systems such as twisted bilayer graphene (TBLG). Strain asymmetrically stretches the hexagonal moiré unit cell of the system, as sketched in Fig. 1a. TBLG samples typically have substantial heterostrain ($\epsilon=0.1\%$$0.4\%$), stemming from several sources during sample preparation and handling, with ultralow-strain samples ($\epsilon<0.1\%$) being rare [nuckollsQuantumTexturesManybody2023, kerelskyMaximizedElectronInterctions2021, turkelOrderlyDisorder2022, kazmierczakStrainFieldsTwisted2021]. Typical and ultralow-strain samples are known to feature different correlated insulating states at low temperatures [nuckollsQuantumTexturesManybody2023], highlighting the need for the modeling of electron correlations in the presence of lattice effects [zhangHeavyFermionsMass2025, caoCorrelatedInsulatorBehaviour2018, nuckollsStronglyCorrelatedChern2021, stepanovCompetingZeroField2021, turkelOrderlyDisorder2022, wongCascadeElectronicTransitions2020, xieSpectroscopicSignaturesManyBody2020, zhangHeavyFermionsMass2025, kazmierczakStrainFieldsTwisted2021].
Lattice relaxation follows from the energetics of the emergent moiré pattern, as locally AB (Bernal) regions are energetically favored compared to the locally AA (aligned) regions. This results in both in-plane and out-of-plane deformations of the atomistic graphene structure, causing the AA regions to shrink relative to the AB regions and bringing the two layers closer to each other at the AB regions compared to the AA regions [caoCorrelatedInsulatorBehaviour2018, choiElectronicCorrelationsTwisted2019, haoElectricFieldTunable2024, xieSpectroscopicSignaturesManyBody2020, yuSpinSkyrmionGaps2023]
@@
Both strain and lattice relaxation effects reduce the symmetries of the system. In the absence of perturbations, TBLG is well described by the seminal Bistritzer-MacDonald (BM) continuum model ([bistritzerMoireBandsTwisted2011] and Fig. 1b) which, in addition to the translational symmetries and $SU(2)$ spin symmetry, possesses $C_{3z},C_{2x},C_{2z},\mathcal{T}$ symmetries and an emergent anti-commuting particle-hole symmetry $P$. Heterostrain generically breaks $C_{3z}$ and $C_{2x}$, splitting the flat band manifold, while lattice relaxation breaks $P$ [herzog-arbeitmanHeavyFermionsEfficient2024], causing the upper flat bands to become more dispersive than the lower flat bands.
In this paper we show that the breaking of some of the lattice and particle-hole symmetries, together with a proper evaluation of electronic correlations, are the key to unlock a unified description of three ubiquitous and hitherto unexplained features observed in TBLG experiments: (i) the emergence of a filling-independent persistent spectral feature at $V_{b}\sim 10$ meV in the local spectral function [wongCascadeElectronicTransitions2020, xiaoInteractingEnergyBands2025, zhangHeavyFermionsMass2025], (ii) a transition from $8$-fold degenerate to $4$-fold degenerate local moments in the flat band manifold when lowering the temperature [zhangHeavyFermionsMass2025, rozenEntropicEvidencePomeranchuk2021, saitoIsospinPomeranchukEffect2021], and (iii) the particle-hole asymmetry of positive vs negative fillings measured from the charge neutrality point (CNP), with charge compressibility varying more strongly on the electron (e)-doped side [saitoIsospinPomeranchukEffect2021, rozenEntropicEvidencePomeranchuk2021, dasSymmetryBrokenChern2021, pierceUnconventionalSequenceCorrelated2021, wongCascadeElectronicTransitions2020], and superconductivity showing higher stability on the hole (h)-doped side [caoUnconventionalSuperconductivityMagic2018, luSuperconductorsOrbitalMagnets2019, yankowitzTuningSuperconductivityTwisted2019] across many samples.
@@
We find that typical strain values $\sim$$0.1\%$ induce a crystal-field–like splitting of the flat-band manifold of the order of $10$ meV. This directly results in a $\sim$$10$ meV persistent feature in the local density, that matches the filling-independent peak reported in [wongCascadeElectronicTransitions2020, xiaoInteractingEnergyBands2025].
Furthermore, our results show that relaxation-induced particle-hole asymmetry in the non-interacting dispersion carries over to the interacting system physics, with stronger variations in the charge compressibility on the electron-doped side compared to the hole-doped side.
@@
Interestingly, we find that in typically strained samples, when the system is doped away from the charge neutrality point, correlations are largely frozen in the inactive (away from the Fermi level) flat-band sector, and present in the active (crossing the Fermi level) flat-band sector. This has strong implications for the degeneracy of the local moments, as is observed in the entropy as as function of filling, which progressively vanishes at the CNP as the temperature is lowered below the characteristic scale of the strain-splitting, in good agreement with recent experimental data [zhangHeavyFermionsMass2025].
@@
In order to correctly describe the dynamical properties of the (magic angle) MA-TBLG electronic structure, we make use of the Topological Heavy Fermion model (THF) [songMagicAngleTwistedBilayer2022], which represents the physics of the flat bands in terms of maximally localized Wannier orbitals accurately mapping the BM continuum model for $1.05^{\circ}$-twisted bilayer graphene to a generalized periodic Anderson model [songMagicAngleTwistedBilayer2022]. Per moiré unit cell, two localized $f$-orbitals (per spin and valley) represent the maximally localized Wannier functions, which are centered in the AA-stacked regions, transform like $p_{x}\pm ip_{y}$ orbitals under the symmetries of TBLG, and capture $\sim 96\%$ of the flat band spectral weight. Four dispersive $c$-orbitals (per spin and valley) make up most of the two closest remote bands above and below the flat-band manifold and guarantee the correct elementary band representation of the flat bands by virtue of mode hybridization at the $\Gamma_{M}$ point. Per spin and valley, the non-interacting Hamiltonian can be written in the form of a $6\times 6$ matrix $H^{\eta}_{HF}(k)$ acting on the spinor $(f^{\eta}_{\alpha},c^{\eta}_{a})$ with $a=1,2,3,4$ and $\alpha=1,2$ representing the orbitals and $\eta=K,K^{\prime}$ representing the valleys.
The THF model can be naturally extended  [herzog-arbeitmanHeavyFermionsEfficient2024, refsupplement] to describe strain and lattice relaxation effects [carrRelaxationDomainFormation2018, vafekContinuumEffectiveHamiltonian2023] which are described by local single particle terms. Ref. [herzog-arbeitmanHeavyFermionsEfficient2024] shows that the modifications to the band structure due to strain and lattice relaxation can be incorporated—directly from ab initio continuum studies [carrRelaxationDomainFormation2018, vafekContinuumEffectiveHamiltonian2023, kangAnalyticalSolutionRelaxed2025]—into the THF model via first-order perturbation theory, resulting in the perturbed Hamiltonian $H^{\eta}_{HF,\epsilon,\Lambda}$,
@@
where $\delta H^{\eta}_{\epsilon}$, $\delta H^{\eta}_{\Lambda}(\mathbf{k})$ represent the corrections due to strain and relaxation respectively, $\epsilon$ parametrizes the uniaxial heterostrain, and $\Lambda$ parametrizes the non-local tunneling terms stemming from lattice relaxation.
@@
The primary effect of $\delta H^{\eta}_{\Lambda}$ is the introduction of two orbital-dependent chemical potential terms that shift the two pairs of $c$-electrons relative to each other and to the $f$-electrons, thus breaking $P$-symmetry. The strain term $\delta H_{\epsilon}^{\eta}$, instead, incites $f$-orbital hybridization causing the flat-band manifold to split. At $0.15\%$ strain, an average value of the typical strain magnitude found in most samples, the split is about $7$ meV. Fig. 1(b) shows the typical band structure with strain and lattice relaxation included. In the following, we will refer to the split narrow bands as the bonding and anti-bonding flat bands, and to the local eigenfunctions of the strained $f$-subspace Hamiltonian (which are eigenstates of $\sigma_{y}$) as the $f_{-}$ and $f_{+}$ orbitals.
To model Coulomb interaction in this setup, we assume a double-gate screened interaction with an inter-gate distance $\xi=10$ nm and dielectric constant $\epsilon=6$. From this, we deduce the value of the many-body interaction coefficients, with the dominant term being the $ff$-interaction $U=57.95$ meV, and the $cf$ and $cc$ terms circa $15$ meV weaker. Given the width of the narrow bands is of the order of $10$ meV [songMagicAngleTwistedBilayer2022, herzog-arbeitmanHeavyFermionsEfficient2024, refsupplement], the interaction-to-bandwidth ratio implies that correlation effects are extremely pronounced. In fact, theoretical simulations show how a mean-field description is not very accurate in capturing spectral and response properties of the system, as well as the temperature scale for phase transitions [raiDynamicalCorrelationsOrder2024, dattaHeavyQuasiparticlesCascades2023]. Hence, we employ the charge self-consistent Dynamical Mean-Field Theory (DMFT) framework developed in [raiDynamicalCorrelationsOrder2024, huSymmetricKondoLattice2023], making use of the CT-QMC solver w2dynamics [wallerbergerW2dynamicsLocalOne2019] to study the effects of correlations in the presence of strain and lattice relaxation in absence of spontaneous symmetry breaking. We treat the $f$-subspace local interactions at all orders, and all other non-local two-body terms, which fall outside the DMFT approximation, at the Hartree level by coupling to the self-consistently adjusted density matrix. We disregard spontaneous symmetry breaking by setting a temperature of $T\geq 11.6K$, above the onset of ordering [raiDynamicalCorrelationsOrder2024]. For more details on the solver and method see [refsupplement].
.3 Spectral persistent feature and charge sector freezing
@@
Figure 2: Momentum-resolved $f$-projected spectral function with lattice relaxation and $0.15\%$ strain at filling $\nu=-0.8$ and $T=11.6K$, projected onto the strain split $f_{+}$ (orange) and $f_{-}$ (blue) orbitals. On the h-doped side ($\nu<0$), the persistent feature at around $10$ meV is made up of unoccupied $f_{+}$ spectral weight.
In Fig. 1(d-f) we show how the spectral function of TBLG evolves with filling when both strain and lattice relaxation are included. The results presented here refer to a strain level $0.15\%$. The dependence of the spectra on the strain mangitude is discussed in the end matter. These figures can be compared directly with the local density measured via scanning tunneling microscopy (STM) [wongCascadeElectronicTransitions2020, kimImagingInterValley2023, nuckollsStronglyCorrelatedChern2021, xieSpectroscopicSignaturesManyBody2020, zhouCoexistenceReconstructedUnreconstructed2023] and the Quantum twisting microscope [inbarQuantumTwistingMicroscope2023, leeRevealingElectronInteractions2025, xiaoInteractingEnergyBands2025], as well as the previous DMFT data on unstrained models [raiDynamicalCorrelationsOrder2024, dattaHeavyQuasiparticlesCascades2023]. Similarly to the unstrained and unrelaxed system, the spectral function features cascade transitions, with side bands forming patterns that repeat every time the number of electrons per moiré unit cell changes by an integer, and the $f$- and $c$-electron occupations resetting. However, the spectra from the perturbed model reproduce one important feature in the experimental data that has thus far eluded explanation. QTM [xiaoInteractingEnergyBands2025] and STM [wongCascadeElectronicTransitions2020] spectra report the presence of a “persistent” (filling-independent) feature at around $\sim$$10$ meV on the h-doped site and $\sim$10-10 meV on the e-doped side. While this feature was absent in previous unstrained DMFT calculations [raiDynamicalCorrelationsOrder2024, dattaHeavyQuasiparticlesCascades2023], our spectral functions of the perturbed model reproduce the persistent feature (cf. the dashed vertical lines in Fig. 1(d-f)) and provide a natural explanation for the corresponding excitations.
Fig. 1(d) shows the local spectral function (i.e. averaged over the mini BZ of TBLG), while (e) and (f) are the spectral functions at the $K_{M}$ and $\Gamma_{M}$ high-symmetry points. The theoretical data are in good agreement with the QTM results at the corresponding momenta [xiaoInteractingEnergyBands2025]. At the $K_{M}$ point, the spectrum presents a well defined gap around the Fermi energy, which is maximum around the CNP. The local spectral function, however, features residual spectral resonance crossing the gap. This is mostly due to the $c$-electron spectral weight at $\Gamma_{M}$, as explained in detail in [refsupplement]. Since the $c$-electrons are weakly correlated, the associated spectral peaks effectively behave as noninteracting bands, rigidly shifting upon chemical potential variations. This gives rise to the striking correspondence between the central spectral peak and the $\mu(\nu)$ curve, represented by the black dotted line in  Fig. 1(f), which is also observed experimentally in [xiaoInteractingEnergyBands2025].
The breaking of $C_{3z}$ by strain causes the $8$-fold degenerate flat-band manifold in the unperturbed THF to split into two sets of degenerate anti-bonding and bonding $f$-bands. In Fig. 1(c), we follow the occupation of the $f_{+}$ and $f_{-}$ electrons as a function of total filling $\nu$. We find that on the e-doped side, the $f_{-}$ manifold occupation is almost complete and constant, while the $f_{+}$ electrons are active, and vice versa on the h-doped side. This suggests that the persistent feature is related to excitations to the inactive sector. Since this switches at the charge neutrality point, it appears at positive or negative bias depending on whether the system is hole- or electron-doped, and the energy at which it appears depends on the magnitude of strain in the sample.
We confirm this hypothesis by looking at the momentum-resolved spectral function in Fig. 2, obtained at $T=11.6K$ and doping $\nu=-0.8$. The hue represents whether the spectral weight comes from the $f_{-}$ (blue) or $f_{+}$ (orange) orbitals. Since this is on the hole-doped side of CNP, the occupation of the $f_{+}$ (orange) electrons is nearly zero. Dynamical correlations are unimportant in this sector and the energy of the corresponding spectral weight is set by the strain splitting. This spectral weight is precisely the content of the persistent feature (marked by the dashed horizontal line in  Fig. 1(e,f,g), and at this filling, is composed of $f_{+}$ (orange) electrons. The active sector is composed of the $f_{-}$ (blue) electrons, whose occupation changes as a function of filling. This sector exhibits dynamical correlations, developing the Hubbard bands and the zero-bias resonance forming the cascades. At the considered filling, this spectral weight is pinned close to the Fermi level.
In agreement with the observations in the QTM experiment [xiaoInteractingEnergyBands2025], this contribution is absent at the $\Gamma_{M}$ point. We arrive at the conclusion that strain causes the flat bands to split (already as a single particle effect) into anti-bonding and bonding subsets, only one of which is active depending on whether the system is electron- or hole-doped. The inactive sector causes the filling-independent persistent feature, while the active sector is responsible for the correlations and cascade resets.
.4 $P$-symmetry breaking and inverse compressibility